OpenStudy (kitkat1):

Devon and his friends bought strawberry wafers for $3 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $30 to buy a total of 22 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of strawberry wafers and the number of packets of chocolate wafers that Devon and his friends bought at the carnival. Define the variables used in the equations. Part B: How many packets of chocolate wafers and strawberry wafers did they buy? Explain how you got the answer and why you selected a

OpenStudy (kitkat1):

WIll give medals and fan!

OpenStudy (anonymous):

What do you think the two variables should represent?

OpenStudy (kitkat1):

idk

OpenStudy (kitkat1):

can u help me with part B

OpenStudy (kitkat1):

Yadu123 can u help me"?

OpenStudy (anonymous):

You need A to do B. What did you get for A?

OpenStudy (kitkat1):

s + c = total number of wafers Part 1. 3 x s + 1 x c = total cost.

OpenStudy (anonymous):

ok. so you need to put the actual numbers from your problem into the equation s + c = 22 3s + c = 30

OpenStudy (kitkat1):

okay what do i do next?

OpenStudy (anonymous):

Subtract the bottom equation from the top to eliminate the variable c.